OpenStudy (anonymous):

Simplify 5^-1/2 * 5x^5/2 (5x)^-3/2 My final answer is x/sqrt5. Is this correct?

OpenStudy (anonymous):

\[5^{-1/2}*5x ^{5/2}(5x)^{-3/2}\]

OpenStudy (anonymous):

Therefore...\[1/\sqrt{5}*\sqrt{5x ^{5}}(1/\sqrt{5x ^{3)}}\]